OpenStudy (anonymous):

write the equation for each parabola with the given conditions: Vertex(-3,-4); the graph passes through the point (1,4). X Intercept of -2 and 1. then find the vertex

OpenStudy (anonymous):

I need all the steps pleased!!!!!!!!!!!

OpenStudy (anonymous):

Is the second line one problem?

OpenStudy (anonymous):

no those are the given conditions you have to find an equation for each parabola write the equation for each parabola with the given conditions: A. Vertex(-3,-4); the graph passes through the point (1,4). B. X Intercept of -2 and 1. then find the vertex

OpenStudy (anonymous):

Oh so they're two different ones. ok well for a, you know the vertex form of a parabola is y=a(x-h)^2+k where (h,k) is the vertex, and since you have another point on the graph, you can plug in the vertex and x and y to solve for a

OpenStudy (anonymous):

SO I KNOW I PLUG IN A(X+3)^2 -4 THEN AFTER THAT WHAT DO I DO?

OpenStudy (anonymous):

Well you have to plug in the other point so you can solve for a.. 4=a(1+3)^2-4

OpenStudy (anonymous):

THEN DO I DISTRIBUTE IT SO IT'LL GIVE ME 4=A(25-4?

OpenStudy (anonymous):

Do the parentheses first 4=16a-4 8=16a 1/2=a So, now that you know that you can say the equation is y=1/2(x+3)^2-4

OpenStudy (anonymous):

THIS IS WHAT I DID 4=A(1+4)^2-4 4=A(5)^2-4 4=A(25)-4 THEN IDK WHAT TO DO

OpenStudy (anonymous):

Well its a 3 inside the parentheses, not a 4

OpenStudy (anonymous):

HOW DID YOU GET 16?

OpenStudy (anonymous):

1+3=4 4^2=16

OpenStudy (anonymous):

ill do part 2 of the problem then

OpenStudy (anonymous):

BUT ITS A -3 NOT A +3?

OpenStudy (anonymous):

The sign changes because its (x-h)

OpenStudy (anonymous):

THEN YOU ONLY DIVIDE 8 BY 16 = 1/2 RIGHT?

OpenStudy (anonymous):

Assuming that this is a normal parabola without any coefficients in front of the x values...... you know your roots, since they are they x-intercepts... with that... you get... \[(x+2)(x-1) = x^{2} + x-2\] which you can find the x-value of the vertex.....

OpenStudy (anonymous):

or you can get your x-value of the vertex since the x-value of the vertex has to be exactly between your roots.... since a parabola is a symmetrical function over it's center x-value or the x-value of the vertex